At a local high school, GPA's are normally distributed with a mean of 2.5 and standard deviation of 0.6. What percentage of students at the high school have a GPA higher than 3.0?



1
Expert's answer
2021-10-17T17:26:12-0400

"P(X>3.0)=P(Z>\frac{3.0-2.5}{0.6})=P(Z>0.83)=1-P(Z<0.83)="

"=1-0.7967=0.2033=20.33\%."